OpenStudy (anonymous):

What is the x-intercept and the y- intercept of the graph of 10x-5y=20?

OpenStudy (mathhelp346):

x intercept=c/a 20/10=2 x intercept=2 y intercept=c/b 20/5=4 y intercept=4 :)

OpenStudy (mathhelp346):

just write down the formulas and you'll get the hang of it slope= (opposite of sign) a/b y intercept= c/b x intercept= c/a

OpenStudy (precal):

Or you can sub x=0 and solve for y (This is your y-intercept) and sub y=0 and solve for x (This is your x-intercept)